OpenStudy (anonymous):

The perimeter of a rectangular field is 348 yards. If the width of the field is 79 yards, what is it lenght?

OpenStudy (anonymous):

\[2(l+w)=348\]

myininaya (myininaya):

Perimeter of a rectangle=2L+2W

myininaya (myininaya):

W=79 Perimeter of rectangle =348

OpenStudy (anonymous):

perimeter = 2(length + width) => 348 = 2(l + 79) => l = 95

OpenStudy (anonymous):

\[l+79=348/2\]

myininaya (myininaya):

348=2L+2(79) you solve for L to find Length

OpenStudy (anonymous):

thanks